OpenStudy (anonymous):

can someone please help me solve this? 5/3 - 3/2x=3/2

OpenStudy (anonymous):

mu;tiply whole equation by 6 it wil be eaier fo ryou then @mlees1969

OpenStudy (anonymous):

(10 x-9)/(6 x) = 3/2

OpenStudy (anonymous):

x = 9

OpenStudy (anonymous):

u can clear ur doubt by seeing this http://www.wolframalpha.com/input/?i=+5%2F3+-+3%2F2x%3D3%2F2

OpenStudy (jiteshmeghwal9):

\[\frac{-3}{2x}=\frac{3}{2}-\frac{5}{3}\]

hero (hero):

0. Original problem \(\large\frac{5}{3} - \frac{3x}{2} =\frac{3}{2}\) 1. Add 3x/2 to both sides: \(\large\frac{5}{3} =\frac{3}{2} + \frac{3x}{2}\) 2. Combine fractions on the right side to get: \(\large\frac{5}{3} =\frac{3 + 3x}{2}\) 3. Cross Multiply to get \(10 = 3(3 + 3x)\) 4. Distribute to get: \(10 = 9 + 9x\) 5. Subract 9 from both sides to get \(1 = 9x\) 6. Divide both sides by 9 to get: \(x = \large\frac{1}{9}\)